OpenStudy (anonymous):

6 fourth root of 9 end root plus 2 cubed root of 9 end root plus 2 fourth root of 9 end root minus 3 cubed root of 9

OpenStudy (anonymous):

?? explain

OpenStudy (jack1):

hiya @carolschroeder so just so i've got this correct, is ur eqn...? \[(6\times \sqrt[4]{9}) + (2\times \sqrt[3]{9}) + (2\times \sqrt[4]{9}) - (3\times \sqrt[3]{9})\]

OpenStudy (jack1):

if so, by adding and subtracting like terms, that would equal \[(8\times \sqrt[4]{9}) - (1\times \sqrt[3]{9})\]

OpenStudy (anonymous):

there cannot be a simplified answer

OpenStudy (jack1):

now as 9 = 3^2 (3 squared) that would be 8 * (3^2)^1/4 - 1* (3^2)^1/3 so that equals 8* 3^(2/4) - 3^(2/3) equals 8* 3^(1/2) - 3^(2/3) so : \[(8\times3^{1/2}) - 3^{2/3} = 8\times \sqrt{3}) - 3^{2/3}\]